OpenStudy (anonymous):

Given that P = (-5, 5) and Q = (-13, 10), find the component form and magnitude of 2. PQ

OpenStudy (anonymous):

Now answer !

OpenStudy (anonymous):

sqrt13??! @E.ali

OpenStudy (anonymous):

It s a formula for you !

OpenStudy (anonymous):

I used the formula and got sqrt13? but I have no clue what I use that for lol...

OpenStudy (anonymous):

@E.ali

ganeshie8 (ganeshie8):

vectors, right ?

OpenStudy (anonymous):

yeah @ganeshie8

OpenStudy (anonymous):

Ok so I did some research of how to get it and I got mag of sqrt89 does that seem right?

ganeshie8 (ganeshie8):

first find the component form, then you need to take mag of component form

OpenStudy (anonymous):

Ok ! To find vectors you should go with another way ...!

OpenStudy (anonymous):

how do I do that? see what I did was I did sqrt(Q1-PI)^2+(Q2-P2)^2

ganeshie8 (ganeshie8):

P = (-5, 5) and Q = (-13, 10) component form : <-13--5, 10-5> <-13+5, 5> <-8, 5>

ganeshie8 (ganeshie8):

thats the component form. to find component form of a vector when its points are given, simply subtract

ganeshie8 (ganeshie8):

now, for mag, take mag of <-8, 5> mag = \(\large \sqrt{x^2+y^2} = \sqrt{(-8)^2+5^2}\) simplify

OpenStudy (anonymous):

see I did that but it wasn't an option so now I am really confused xD LOL

OpenStudy (anonymous):

since it says mag 2 should I multiply <-8,5> by two?

ganeshie8 (ganeshie8):

Yes ! after taking the mag, multiply it by 2

OpenStudy (anonymous):

so it would be sqrt178 and <-16,10>

OpenStudy (anonymous):

:DDD!?

ganeshie8 (ganeshie8):

2PQ = \(\large 2 \times \sqrt{(-8)^2 + 5^2 }\)

OpenStudy (anonymous):

is that right?? :D

ganeshie8 (ganeshie8):

that should give you 2 x sqrt(89) = 18.87

ganeshie8 (ganeshie8):

is that one of your options ha ? :)

OpenStudy (anonymous):

Yeah it is!!!! !:DDDDD THANK YOU SOOOOO MUCH! !<3

ganeshie8 (ganeshie8):

wonderful !!! you wlcme :)